Project

General

Profile

02042015

Attendance

Louise, Fernanda, Ryan, Justin, Leon, Ash River.

Notes

FD good runs are showing only preliminary data on day plot.
  • Action: email Joao (done). there was an issue with some subruns not being processed, nearline experts are now aware.

FD noise:
Noticed on week plots number of noise hits had decreased. Nothing to suggest why in log book.
On month plots also see bigger decreases in noise but they correspond to work on detector.

FD slice standard dev.
Odd spikes in plot. No one had an idea as cause. Leon says other triggers but these plots should just be numu, yes?

ND noise:
Looks like its stabilizing? Only since the 31st so hard to judge.

There was dissuasion of metrics to map the number of hardware issues we have over time on a channel by channel basis if possible.
Coincidentally Keith gave a talk showing similar plots in Opps meeting.
  • Action. Email tread started with him, Fernanda, Ryan discussing this.

Gave a brief update at opps meeting.
Generally talked to go down well at operations meeting, people seemed interested.

  • Action. Jaroslav asked if we could get zoomed versions of FEB drop out and noise rate vs time plots.